### v1

# Partial Resolution of the Erdős-Straus, Sierpiński, and Generalized Erdős-Straus Conjectures Using New Analytical Formulas

Abstract
This article proposes a unified analytical approach leading to a partial resolution of the Erdős–Straus, Sierpiński conjectures, and their generalization. We introduce two new analytical formulas under specific conditions of divisibility and the existence of perfect squares. Under these conditions, the formulas verify the conjectures even for very large numerical values. Moreover, our method reduces the problem to the search for a suitable perfect square, thereby opening the way to a complete proof of these conjectures. Notably, our second formula significantly improves upon Mordell’s work by demonstrating analytically the conjecture in the majority of cases where Mordell’s approach fails. Furthermore, these formulas are highly versatile, as they provide, under the established conditions, a systematic method to decompose any fraction $\frac{a}{n}$ into a sum of three Egyptian fractions. In conclusion, we present open questions and conjectures to the mathematical community regarding the generalization of these formulas.
1 Introduction
The Erdős-Straus conjecture, formulated in 1948 by Paul Erdős and Ernst G. Strauss, states that for every natural number $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that:
$\frac{4}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$ (1)
Louis Mordell demonstrated that if $n$ is not congruent to $1,11^{2},13^{2},17^{2},19^{2},23^{2}$ modulo 840, then the conjecture holds. Furthermore, numerical verifications have confirmed its validity up to $n=10^{17}$.
For decades, mathematicians have been intrigued by the following question: Does an explicit formula exist relating $x,y,z$, and $n$ that systematically verifies the conjecture?
In this article, we propose two new analytical formulas obtained through an in-depth study of the Diophantine equation (1). The first relies on a natural divisibility condition, while the second is based on the existence of a perfect square. These results provide a partial proof of the conjecture and represent a significant advancement in its study.
The Polish mathematician Wacław Sierpiński also investigated this conjecture and posed a similar question: Is a similar decomposition possible if 4 is replaced by 5? He conjectured that for every natural number $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that:
$\frac{5}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$ (2)
In this article, we will demonstrate that our formulas also verify this conjecture under the same conditions as before.
Finally, following the formulation of the Erdős-Straus conjecture, several mathematicians explored the possibility of generalizing this Egyptian fraction representation. This led to the generalized Erdős-Straus conjecture, which states that for every $a\geq 4$ and every $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that:
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$ (3)
We will show that our two analytical formulas also verify this conjecture.
In this article, we present a detailed analytical approach leading to these formulas, along with numerical tests supporting our results.
2 The Erdős-Straus Conjecture
One of the methods that is known nowadays is that if $n$ is even integer , meaning $n=2k$ with $k\in\mathbb{N}^{*}$. In this case, it is well established that:
$\frac{4}{n}=\frac{1}{\frac{n}{2}}+\frac{1}{n}+\frac{1}{n}.$ (4)
This result is one of the most trivial among those obtained since the conjecture was stated, as it shows that all even numbers satisfy the conjecture.
In my new approach, I will demonstrate how this result can be naturally recovered when $n$ is even. Before introducing the general method that allowed me to establish two formulas verifying the conjecture under a certain condition, I will first illustrate my process by analytically testing the conjecture for some values of $n$.
My method is based on the structural analysis of equation (1). It is a simple yet powerful and general approach, applicable to even numbers, odd numbers, and prime numbers alike.
2.1 Case n = 2 (Even Number)
The fundamental equation becomes:
$\frac{4}{2}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$
I choose not to simplify $\frac{4}{2}$ immediately in order to maintain the general structure of the method. However, one can simplify this fraction directly, and by following the same steps that I will present, the same result is always obtained.
The equation can then be rewritten as:
$\frac{4}{2}-\frac{1}{x}=\frac{1}{y}+\frac{1}{z},$
which reformulates as:
$\frac{4x-2}{2x}=\frac{y+z}{yz}.$
Since we are searching for solutions in $\mathbb{N}^{*}$, it is necessary that $24x-$ 2 be strictly positive, that is, $04x-2>$ 0, which gives $x\geq\lfloor\frac{1}{2}\rfloor+1=1$.
For the rest of the document, the notation $\lfloor x\rfloor$ will always denote the integer part of $x$ to avoid any confusion. We then obtain the equation:
$yz(4x-2)=2x(y+z).$
At this stage, I apply a method commonly used in geometry of curves and surfaces, called parameterization.
2.2 Parameterization Technique
Let $a,b,c,d\in\mathbb{R}$ such that:
$ab=cd.$
Then, for any $t\in\mathbb{R}$, we can set:
$a=dt,\quad c=bt.$
This parametric representation holds for any $t$ and allows rewriting $(dt)b=(bt)d$, which verifies the initial equation. This technique will be crucial for my results.
2.3 Application to Our Equation
We then set:
$y+z=(4x-2)t,\quad yz=2xt.$
This parameterization is valid for all $t\in\mathbb{N}^{*}$, as we are looking for solutions in $\mathbb{N}^{*}$.
According to the sum and product method, **$y$ and $z$ are solutions of the quadratic equation:**
$V^{2}-((4x-2)t)V+2xt=0.$
Its discriminant is given by:
$\Delta=t^{2}(4x-2)^{2}-8xt.$
We imposed $x\geq 1$. Choosing $x=1$, we obtain:
$\Delta=4t^{2}-8t=t(4t-8).$
For $y$ and $z$ to be integers, $\Delta$ must be a perfect square, and $24x-$ 2 must have the same parity as $\sqrt{\Delta}$.
Taking $t=2$ (a trivial value for which $\Delta=0$), we obtain a double root:
$y=z=\frac{(4\times 1-2)\times 2}{2}=2.$
Thus, for $n=2$, we find $x=1,y=2,z=2$, verifying:
$\frac{4}{2}=\frac{1}{1}+\frac{1}{2}+\frac{1}{2}.$
This confirms the validity of the conjecture for $n=2$.
2.4 Analysis of the Method
if we had directly applied the known formula for even numbers given by equation ($4$) , we would have found:
$x=\frac{2}{2}=1,\quad y=z=2.$
Thus, we obtain exactly the same values.
One of the key advantages of our analytical approach is its applicability not only to even numbers but also to odd numbers and prime numbers. Unlike classical methods-which have relied primarily on brute-force techniques that merely test the existence of solutions without offering insight into the problem’s structure—our method derives solutions directly from the inherent logical structure of the underlying Diophantine equation. This provides a general, rigorous framework for addressing the Erdős-Straus, Sierpiński, and generalized Erdős-Straus conjectures. Moreover, the elegance and clarity of our approach make it a valuable pedagogical tool for teaching methods of solving Diophantine equations, as it not only yields the solutions but also deepens our understanding of the arithmetic properties involved
Of course, testing the conjecture for specific values of $n$ does not constitute a general proof. My goal here is to demonstrate the efficiency of this method in solving equation (1) analytically.
Next, we will examine the cases $n=3$, $n=17$, and finally $n=841$ (one of the numbers that Mordell had not proven), before moving on to generalization. Since the main challenge concerns odd numbers, I will not dwell much on even numbers. However, you can reproduce this reasoning for any even $n$, and it will work the same way.
2.5 Case $n=3$ (odd and prime Number)
Applying the same procedure, we impose:
$x\geq\left\lfloor\frac{3}{4}\right\rfloor+1=1.$
We then obtain the system:
$\begin{cases}y+z=(4x-3)t,\\ yz=3xt,\end{cases}$
for all $t\in\mathbb{N}^{*}$.
The values of $y$ and $z$ satisfy the quadratic equation:
$V^{2}-((4x-3)t)V+3xt=0.$
The discriminant is given by:
$\Delta=t^{2}(4x-3)^{2}-12xt.$
We choose $x\geq\left\lfloor\frac{4}{3}\right\rfloor+1=2$.
Taking $x=2$, we obtain:
$\Delta=25t^{2}-24t=t(25t-24).$
A trivial solution ensuring that $\Delta$ is a perfect square is $t=1$, yielding $\Delta=1$.
Thus, we find:
$y=\frac{(4\times 2-3)-1}{2}=2,\quad z=\frac{(4\times 2-3)+1}{2}=3.$
For $n=3$, we have $x=2$, $y=2$, and $z=3$, verifying:
$\frac{4}{3}=\frac{1}{2}+\frac{1}{2}+\frac{1}{3}.$
By selecting a different $x$, we can determine corresponding values of $t$ that yield integer solutions. For instance, choosing $x=4$ results in:
$\Delta=169t^{2}-48t=t(169t-48).$
For $t=1$, we find a perfect square $12111^{2}=$ 121, yielding $y=1$ and $z=12$.
2.6 Case $n=17$ (odd and prime Number)
We impose:
$x\geq\left\lfloor\frac{17}{4}\right\rfloor+1=5.$
The system obtained is:
$\begin{cases}y+z=(4x-17)t,\\ yz=17xt,\end{cases}$
for all $t\in\mathbb{N}^{*}$.
$y$ and $z$ satisfy the quadratic equation:
$V^{2}-((4x-17)t)V+17xt=0.$
The discriminant is:
$\Delta=t^{2}(4x-17)^{2}-68xt.$
To minimize $(4x-17)^{2}$ and simplify calculations, we choose the smallest $x\geq 5$, i.e., $x=5$.
This gives:
$\Delta=9t^{2}-340t=t(9t-340).$
We require $\Delta$ to be a perfect square, meaning $9t-340\geq 0$, so $t\geq\frac{340}{9}\approx 37.77$.
Using a Python script to scan $t$ in the range $[38,1000]$, we find the perfect square values of $\Delta$:
$t=68\Rightarrow\sqrt{\Delta}=136,\quad t=180\Rightarrow\sqrt{\Delta}=480.$
For $t=68$, we obtain:
$y=34,\quad z=170.$
Thus, we confirm:
$\frac{4}{17}=\frac{1}{5}+\frac{1}{34}+\frac{1}{170}.$
The reader may compute $y$ and $z$ for $t=180$ using the same approach.
2.7 Case $n=841$ (number not proven by Mordell)
Following the same procedure, we impose:
$x\geq\left\lfloor\frac{841}{4}\right\rfloor+1=211.$
The system obtained is:
$\begin{cases}y+z=(4x-841)t,\\ yz=841xt,\end{cases}$
for all $t\in\mathbb{N}^{*}$.
$y$ and $z$ satisfy the quadratic equation:
$V^{2}-((4x-841)t)V+841xt=0.$
The discriminant is:
$\Delta=t^{2}(4x-841)^{2}-3364xt.$
Choosing $x=211$, we obtain:
$\Delta=9t^{2}-709804t=t(9t-709804).$
For $\Delta$ to be a perfect square, we require $9t-709804\geq 0$, so $t\geq\frac{709804}{9}\approx 78867.11$.
Scanning for $t$ in the range $[78868,500000]$, we find:
$t=185600,\quad\sqrt{\Delta}=422240.$
Thus, we compute:
$y=67280,\quad z=489520.$
Verifying:
$\frac{4}{841}=\frac{1}{211}+\frac{1}{67280}+\frac{1}{489520}.$
Both sides yield the same numerical value.
3 Conclusion
This analytical approach effectively solves equation (1) and finds solutions satisfying the conjecture for even, odd, and prime numbers.
With this method, one can manually prove the conjecture for many numbers not previously proven by Mordell.
However, verifying the conjecture for specific values does not constitute a general proof. We will now proceed to the generalization and present the results obtained.
4 Generalization of the Erdős-Straus Conjecture
We want to prove that for all $n$ in $\mathbb{N}^{*}$ with $n\geq 2$, there exist $x,y,z$ in $\mathbb{N}^{*}$ such that
$\frac{4}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}$
We have called this equation (1).
My proposed proof follows the same structure as before, as you may have noticed, the value of $n$ does not change the structure of the proof. I will use exactly the same procedure as previously.
Equation (1) is equivalent to:
$\frac{4}{n}-\frac{1}{x}=\frac{y+z}{yz}$
which is also equivalent to
$\frac{4x-n}{nx}=\frac{y+z}{yz}$
For the right-hand side to be strictly positive, it is necessary that $04x-n>$ 0, i.e., $x>\frac{n}{4}$, so we have $x\geq\left\lceil\frac{n}{4}\right\rceil$.
We still have:
$\frac{4x-n}{nx}=\frac{y+z}{yz}$
which holds if and only if
$yz(4x-n)=nx(y+z)$
Similarly, using parameterization, we can express:
$\begin{cases}y+z=(4x-n)(2t)\\ yz=nx(2t)\end{cases}$
for all $t$ in $\mathbb{N}^{*}$. If we replace $y+z$ and $yz$ with these expressions, the equation $yz(4x-n)=nx(y+z)$ will always hold.
Later, I will explain why I parameterized using $2t$ in this general case instead of using $t$ as I did for specific values of $n$. (Of course, we are allowed to parameterize using any multiple of $t$ in $\mathbb{N}^{*}$, as the equation will always hold.)
Now, let us use sums and products. From the system above, $y$ and $z$ satisfy the quadratic equation in $V$:
$V^{2}-(\text{sum})V+\text{product}=0$
which gives:
$V^{2}-((4x-n)(2t))V+2ntx=0.$
The discriminant is given by:
$\Delta=4t^{2}(4x-n)^{2}-8ntx.\quad(5)$
You may recall that I promised to retrieve equation (4) for even values of $n$. If we set $n=2x$, meaning $n$ is even since $x$ is in $\mathbb{N}^{*}$, the discriminant becomes:
$\Delta=4t^{2}(4x-2x)^{2}-16x^{2}t.$
$=16x^{2}t^{2}-16x^{2}t.$
An obvious choice for $t$ so that $\Delta$ is a perfect square is $t=1$, which gives $\Delta=0$.
Thus,
$y=z=\frac{2(4x-2x)}{2}=2x$
so $y=z=n=2x$, which implies $x=\frac{n}{2}$, and we obtain:
$\frac{4}{n}=\frac{1}{\frac{n}{2}}+\frac{1}{n}+\frac{1}{n}$
which retrieves equation (4).
This was a side note, as our goal is to generalize. So, we return to our quadratic equation in $V$ in its general form, with $n\in\mathbb{N}^{*}$ such that $n\geq 2$ (as required by the conjecture).
We still have our discriminant computed in (5):
$\Delta=4t^{2}(4x-n)^{2}-8ntx.$
Expanding this expression gives:
$\Delta=64t^{2}x^{2}-32t^{2}xn+4t^{2}n^{2}-8ntx.\quad(6)$
Now, let us determine for which values of $x$ we have $\Delta>0$. From (6), solving $\Delta>0$ is equivalent to solving:
$t(64tx^{2}-32txn+4tn^{2}-8nx)>0$
Since $t>0$, this simplifies to solving:
$064tx^{2}-32txn+4tn^{2}-8nx>$ 0 .
Solving this inequality gives:
$x\in\left]0,\frac{n(4t+1-\sqrt{8t+1})}{16t}\right[\cup\left]\frac{n(4t+1+\sqrt{8t+1})}{16t},+\infty\right[$
where $x$ belongs to $\mathbb{N}^{*}$.
To obtain integer solutions for this equation, it is necessary that $\Delta$ be a perfect square. I struggled to find an obvious perfect square for $\Delta$. It seems to me that the polynomial given in (6) is rich, but unfortunately, after some quick and unsuccessful attempts to find a classic perfect square of $\Delta$ without imposing any condition on $n$ (since we are in the process of generalization), the idea I had was to use the fact that the discriminant $\Delta$ holds for any $t\in\mathbb{N}^{*}$ and see what can be deduced to obtain a perfect square in order to bypass the difficulty. Let’s see how I proceed Let’s revisit the compact form of $\Delta$ in (5).
5 Compact Form of $\Delta$
$\Delta=4t^{2}(4x-n)^{2}-8ntx$
This simplifies to:
$\Delta=4t(t(4x-n)^{2}-2nx).$
This is valid for all $t\in\mathbb{N}^{*}$ because, as mentioned earlier, the parametrization is valid for all $t\in\mathbb{N}^{*}$. You also observed how this information allowed us to prove the conjecture (manually for some $n$) by providing specific values of $t$ that help us obtain a perfect square. We will use this information in this generalization.
We saw that $\Delta>0$ for
$x\in\left]0,\frac{n(4t+1-\sqrt{8t+1})}{16t}\right[\cup\left]\frac{n(4t+1+\sqrt{8t+1})}{16t},+\infty\right[$
Recalling the expression for $\Delta$,
$\Delta=4t(t(4x-n)^{2}-2nx),$
and since $\Delta>0$ for $x\in\left]0,\frac{n(4t+1-\sqrt{8t+1})}{16t}\right[\cup\left]\frac{n(4t+1+\sqrt{8t+1})}{16t},+\infty\right[$, for the x and t which respect this condition, it is clear that there exist $x$ and $t$ such that $t(4x-n)^{2}-2nx$ is in $\mathbb{N}^{*}$. We can then write:
$t(4x-n)^{2}-2nx=k,$
with $k\in\mathbb{N}^{*}$ Thus,
$\Delta=4t(t(4x-n)^{2}-2nx)=4tk,$
which holds for all $t\in\mathbb{N}^{*}$. In particular, for $t=k$ (This is why we excluded the case $\Delta=0$, otherwise $k=\left(t(4x-n)^{2}-2nx\right)$ can be zero and we cannot set $t=k$ because from the start, during the parameterization, $t$ is non-zero). , so we have:
$\Delta=4tk=4k^{2}=(2k)^{2}.$
We have thus obtained a perfect square. Let’s now see under what conditions this transformation holds.
6 Conditions for Obtaining a Perfect Square
To obtain this perfect square, we wrote $t(4x-n)^{2}-2nx=k$, and then set $t=k$, simplifying the equation to:
$k(4x-n)^{2}-2nx=k,$
or:
$k\left((4x-n)^{2}-1\right)=2nx.$
This implies that:
$k=\frac{2nx}{(4x-n-1)(4x-n+1)},$
and $k\in\mathbb{N}^{*}$ if and only if $(4x-n-1)(4x-n+1)$ divides $2nx$. This means that both $(4x-n-1)$ divides $2nx$ and $(4x-n+1)$ divides $2nx$.
Under these conditions, $\Delta=4k^{2}=(2k)^{2}$, so:
$V=\frac{(4x-n)(2k)\pm 2k}{2}.$
there is a simplification of 2, This is why I had previously parameterized by $2t$, and we just replaced $t$ with $k$ to calculate the roots of this polynomial, and this was possible because we set $t=k$ to obtain a perfect square.
we have $y=k(4x-n-1)$ and $z=k(4x-n+1)$. Now, from earlier, we know that:
$k=\frac{2nx}{(4x-n-1)(4x-n+1)}$
Substituting $k$ into the expressions for $y$ and $z$, we get:
$y=\frac{2nx}{(4x-n-1)(4x-n+1)}(4x-n-1)=\frac{2nx}{4x-n+1}$
and
$z=\frac{2nx}{(4x-n-1)(4x-n+1)}(4x-n+1)=\frac{2nx}{4x-n-1}$
Next, we find the conditions for $x$ such that both $y$ and $z$ are positive. Since $02nx>$ 0, we only need to ensure that the denominators are strictly positive.
Case for $y$:
$y>0$ if:
$4x-n+1>0\quad\text{which implies}\quad x>\frac{n-1}{4}$
Thus, for $y$, we need $x\geq\left\lceil\frac{n-1}{4}\right\rceil+1$.
Case for $z$:
$z>0$ if:
$4x-n-1>0\quad\text{which implies}\quad x>\frac{n+1}{4}$
Thus, for $z$, we need $x\geq\left\lceil\frac{n+1}{4}\right\rceil+1$.
Therefore, the overall condition is:
$x\geq\max\left\{\left\lceil\frac{n-1}{4}\right\rceil+1,\left\lceil\frac{n+1}{4}\right\rceil+1\right\}$
which gives:
$x\geq\left\lceil\frac{n+1}{4}\right\rceil+1$
Thus, $x\in\left[\left\lceil\frac{n+1}{4}\right\rceil+1,+\infty\right[$.
Now, let us check for which values of $x$, $y$, and $z$ are in $\mathbb{N}^{*}$. We recall the following property: if $a\in\mathbb{N}^{*}$ and $b\in\mathbb{N}$, and if $a$ divides $b$, then $a\leq b$.
Therefore, $y\in\mathbb{N}^{*}$ if and only if $14x-n+$ 1 divides $2nx$, which implies:
$4x-n+1\leq 2nx$
$14x-2nx\leq n-$ 1
$x(4-2n)\leq n-1$
For $n>2$, $04-2n<$ 0, so:
$x\geq\frac{n-1}{2n-4},\quad x\in\mathbb{N}$
Similarly, we can show that $z\in\mathbb{N}^{*}$ if:
$x\geq\frac{n+1}{2n-4}$
Thus, we have the following conditions:
$x\geq\left\lceil\frac{n+1}{4}\right\rceil+1$
$x\geq\frac{n-1}{2n-4}$
$x\geq\frac{n+1}{2n-4}$
The two latter values of $x$ are not valid for $n=2$. Therefore, the correct choice that satisfies all conditions is:
$x\geq\left\lceil\frac{n+1}{4}\right\rceil+1$
In other words, $x\in\left[\left\lceil\frac{n+1}{4}\right\rceil+1,+\infty\right[$.
As we will see in the numerical checks, under these divisibility conditions, this formula will allow us to decompose the fraction $\frac{4}{n}$ into three unit fractions for those $n$ satisfying these conditions up to infinity (since we have shown that the values of $x$ for which $14x-n-$ 1 and $14x-n+$ 1 divide $2nx$ extend to infinity). This is good news because we always hope that for any given $n$, there is an infinite number of choices for $x$ that satisfy the divisibility condition, allowing us to prove the conjecture for that specific $n$.
Before proceeding with this, let’s first verify whether, under these divisibility conditions, the Erdős–Strauss conjecture is satisfied for:
$y=\frac{2nx}{4x-n+1}\quad\text{and}\quad z=\frac{2nx}{4x-n-1}$
We need to check if:
$\frac{1}{x}+\frac{1}{y}+\frac{1}{z}=\frac{4}{n}$
Substituting the expressions for $y$ and $z$:
$\frac{1}{x}+\frac{1}{\frac{2nx}{4x-n+1}}+\frac{1}{\frac{2nx}{4x-n-1}}=\frac{1}{x}+\frac{4x-n+1}{2nx}+\frac{4x-n-1}{2nx}$
Simplifying:
$\frac{1}{x}+\frac{(4x-n+1)+(4x-n-1)}{2nx}=\frac{1}{x}+\frac{8x-2n}{2nx}=\frac{1}{x}+\frac{4x-n}{nx}=\frac{4}{n}$
Thus, $x$, $y$, and $z$ satisfy the Erdős–Strauss conjecture under the established divisibility conditions.
In summary, we have proven that for all $n\geq 2$, If there exists $x_{0}\in\mathbb{N}^{*}$ with $x_{0}\in\left[\left\lceil\frac{n+1}{4}\right\rceil+1,+\infty\right[$. such that $14x_{0}-n-$ 1 and $14x_{0}-n+$ 1 divide $2nx_{0}$, then there exist $y=\frac{2nx_{0}}{4x_{0}-n-1}$ and $z=\frac{2nx_{0}}{4x_{0}-n+1}$ in $\mathbb{N}^{*}$ such that:
$\frac{4}{n}=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z}$
Below, I will provide numerical calculations that, depending on the range searched for $x_{0}$, cover a some $n$ Based on this formula verifying the conjecture, you will also be able to perform calculations using more powerful computational resources to identify very large $x_{0}$ and $n$. Of course, with numerical computations, it is impossible to reach +$\infty$+ ∞, but what assures us of the validity of the results is the analytical result we have just established. Every time we find such an $x_{0}$, the conjecture holds for the corresponding $n$.
My question to the mathematical community for a complete proof of the conjecture is as follows: How can we explicitly find $x_{0}$? If it is not possible to obtain an explicit expression for $x_{0}$ (which may depend on $n$), can we prove the existence of $x_{0}$ by induction on $n$ (whether classical or strong)? I remain available for any suggestions or collaborations to address this question and definitively resolve the conjecture.
7 Second Formula: Condition of Existence of a Perfect Square
The first formula above is based on a divisibility condition, while the second one, which I will now introduce, relies on the existence of a perfect square. Let us reconsider the discriminant $\Delta$ of our quadratic equation:
$\Delta=4t^{2}(4x-n)^{2}-8ntx=4(t^{2}(4x-n)^{2}-2ntx).$
We have already seen that $\Delta\geq 0$ for $x\in\left]0,\frac{n\left(4t+1-\sqrt{8t+1}\right)}{16t}\right]\cup\left[\frac{n\left(4t+1+\sqrt{8t+1}\right)}{16t},+\infty\right[$ with $x\in\mathbb{N}^{*}$. Therefore, for $x$ that satisfies this condition, we define
$t^{2}(4x-n)^{2}-2ntx=q^{2},\quad q\in\mathbb{N}.$
Thus, we have:
$\Delta=4q^{2}=(2q)^{2}.$
We will now prove that the conjecture is true under the condition that this perfect square exists.
Let $y=t(4x-n)-q$ and $z=t(4x-n)+q$. We need to show that, with this formula, the conjecture holds for those $q$ such that
$t^{2}(4x-n)^{2}-2ntx=q^{2},\quad q\in\mathbb{N}.$
It is sufficient to check if
$\frac{1}{x}+\frac{1}{y}+\frac{1}{z}=\frac{1}{x}+\frac{1}{t(4x-n)-q}+\frac{1}{t(4x-n)+q}.$
We can simplify this expression:
$\frac{1}{x}+\frac{1}{t(4x-n)-q}+\frac{1}{t(4x-n)+q}=\frac{1}{x}+\frac{2t(4x-n)}{t^{2}(4x-n)^{2}-q^{2}}.$
Now, since $q^{2}=t^{2}(4x-n)^{2}-2ntx$, we have:
$\frac{1}{x}+\frac{2t(4x-n)}{t^{2}(4x-n)^{2}-t^{2}(4x-n)^{2}+2ntx}=\frac{1}{x}+\frac{2t(4x-n)}{2nxt}.$
Simplifying further:
$=\frac{1}{x}+\frac{4x-n}{nx}=\frac{n-n+4x}{nx}=\frac{4}{n}.$
Thus, the second formula allows us to affirm that for any $n\in\mathbb{N}$ with $n\geq 2$, if there exists $(x_{0},t_{0})\in(\mathbb{N}^{*})^{2}$ such that
$t_{0}^{2}(4x_{0}-n)^{2}-2nt_{0}x_{0}=q^{2}\in\mathbb{N},$
then there exist $y=t_{0}(4x_{0}-n)-q$ and $z=t_{0}(4x_{0}-n)+q$ with $(y,z)\in\mathbb{N}^{2}$, such that
$\frac{4}{n}=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z}.$
This also proves the conjecture under this condition. I will also present below the numerical values of $n$ found within a range where we search for $(x_{0},t_{0})$ through numerical calculations.
I also pose the following question to the mathematical community: how can we find an explicit formula for $(x_{0},t_{0})$ that works for all $n$ without any conditions? In other words, how can we prove that the polynomial $t^{2}(4x-n)^{2}-2ntx$ or $4t^{2}(4x-n)^{2}-8ntx$ always admits a perfect square for all $n\in\mathbb{N}$, with $n\geq 2$, in order to complete the proof of the conjecture?
I remain available for suggestions or collaborations to address this question.
This is the contribution of this article to the Erdős-Strauss conjecture.
8 Sierpiński Conjecture
The Sierpiński conjecture is similar to the Erdős-Straus conjecture, except that 4 is replaced by 5. It states that for any $n\in\mathbb{N}$ with $n\geq 2$, there exist $x,y,z\in\mathbb{N}^{*}$ such that:
$\frac{5}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$ (7)
A trivial case arises when $n$ is divisible by 3. In this case, one can simply take $x=\frac{n}{3}$, $y=z=n$, and the conjecture is verified for all $n$ divisible by 3.
The good news is that with this new approach, everything stated for the Erdős-Straus conjecture remains valid for the Sierpiński conjecture by simply replacing 4 with 5. This demonstrates the power and generality of this method.
To keep the article concise, I will not manually test specific values of $n$ as I did for the Erdős-Straus conjecture. However, the reasoning remains the same, and the reader may verify this independently.
8.1 Transforming the Equation
Equation (7) is equivalent to:
$\frac{5}{n}-\frac{1}{x}=\frac{y+z}{yz},$ (4)
which can be rewritten as:
$\frac{5x-n}{nx}=\frac{y+z}{yz}.$ (5)
For the right-hand side to be strictly positive, it is necessary that $05x-n>$ 0, i.e., $x>\frac{n}{5}$, leading to $x\geq\left\lfloor\frac{n}{5}\right\rfloor+1$.
Using parameterization, we set:
$y+z=(5x-n)(2t),\quad yz=nx(2t),\quad\forall t\in\mathbb{N}^{*}.$ (6)
From the above, $y$ and $z$ satisfy the quadratic equation:
$V^{2}-(5x-n)(2t)V+2ntx=0.$ (7)
The discriminant of this equation is:
$\Delta=4t^{2}(5x-n)^{2}-8ntx.$ (8)
8.2 Special Case: $n$ Divisible by 3
If $n=3x$, then:
$\Delta$ $=4t^{2}(5x-3x)^{2}-24x^{2}t,$
$=16x^{2}t^{2}-24x^{2}t.$
A simple choice of $t$ making $\Delta$ a perfect square is $t=2$, yielding $\Delta=16x^{2}=(4x)^{2}$.
Thus, solving for $y$ and $z$:
$y$ $=\frac{4(5x-3x)-4x}{2}=2x=\frac{2n}{3},$
$z$ $=2n.$
Hence, we confirm:
$\frac{5}{n}=\frac{1}{n/3}+\frac{1}{2n/3}+\frac{1}{2n}.$ (9)
This case was just an example; our goal is to generalize the approach.
8.3 General Case: Finding $x$ Values for which $\Delta>0$
Expanding $\Delta$:
$\Delta=100t^{2}x^{2}-40t^{2}xn+4t^{2}n^{2}-8ntx.$ (10)
Solving for $x$ such that $\Delta>0$, we obtain:
$x\in\left]0,\frac{n(5t+1-\sqrt{10t+1})}{25t}\right[\cup\left]\frac{n(5t+1+\sqrt{10t+1})}{25t},+\infty\right[.$ (11)
Since:
$\Delta=4t^{2}(5x-n)^{2}-8ntx=4t(t(5x-n)^{2}-2nx),$ (12)
we can set $t=k=k(5x-n)^{2}-2nx$, ensuring that $k\in\mathbb{N}^{*}$ if $15x-n-$ 1 and $15x-n+$ 1 divide $2nx$.
Thus, we find:
$\Delta=4k^{2}=(2k)^{2},$ (13)
leading to:
$y=k(5x-n-1),\quad z=k(5x-n+1).$ (14)
Replacing $k$ by its expression:
$k=\frac{2nx}{(5x-n-1)(5x-n+1)},$ (15)
we obtain:
$y$ $=\frac{2nx}{5x-n+1},$
$z$ $=\frac{2nx}{5x-n-1}.$
It follows that:
$\frac{5}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$ (16)
8.4 Conclusion
All details provided for the Erdős-Straus conjecture remain valid here, so I will not repeat them. I will also include numerical computations generating certain $n$ values that satisfy these conditions.
Similarly, for the second approach:
$\Delta=4(t^{2}(5x-n)^{2}-2ntx),$ (17)
for values of $x$ satisfying $\Delta>0$, we set
$t^{2}(5x-n)^{2}-2ntx=q^{2},\quad q\in\mathbb{N}.$
, leading to:
$\Delta=4q^{2}=(2q)^{2}.$ (18)
We will now prove that the conjecture is true under the condition that this perfect square exists.
We have:
$y=\frac{2t(5x-n)-2q}{2}=t(5x-n)-q,\quad z=t(5x-n)+q.$
Similarly, we can see that:
$\frac{5}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$
By substituting $q^{2}$ with $t^{2}(5x-n)^{2}-2ntx$, we obtain the desired result.
To conclude, for any $n\geq 2$, if there exists $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that:
$t_{0}^{2}(5x_{0}-n)^{2}-2nt_{0}x_{0}=q^{2}\in\mathbb{N},$ (19)
then $y=t_{0}(5x_{0}-n)-q$ and $z=t_{0}(5x_{0}-n)+q$ satisfy the equation, proving the conjecture under this condition.
I invite the mathematical community to find a general explicit formula for $(x_{0},t_{0})$ that works for all $n$, thereby fully resolving the Sierpiński conjecture.
9 Generalization of the Erdős-Straus and Sierpiński Conjectures
We recall that the generalized of the Erdős-Straus and Sierpiński Conjectures states that for any nonzero natural number $a$ with $a\geq 4$, and for any natural number $n$, $n\geq 2$, there exist $x,y,z$ in $\mathbb{N}^{*}$ such that:
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$
It should be noted that since the inception of this conjecture, there has not been significant research on it, and thus no concrete leads have emerged. In this paper, I will attempt to partially prove it as I did with the two conjectures above, and I will reduce the difficulty to verifying two conditions to completely settle all three conjectures. Here, I will use the same approach as in the previous two conjectures.
It is evident that the Erdős-Straus conjecture and the Sierpiński conjecture are special cases of the generalized Erdős-Straus and Sierpiński conjecture. Therefore, all results established here remain valid for the two conjectures discussed earlier.
9.1 Trivial Case
As before, let us start with the trivial case: If $(a-2)$ divides $n$, then we can simply take $x=\frac{n}{a-2}$ and $y=z=n$, which proves the conjecture in this particular case.
9.2 Equivalence of the conjecture
We will show that the generalized Erdős-Straus and Sierpiński conjecture is true if and only if $y$ and $z$ are solutions of the following quadratic equation in $V$:
$V^{2}-(ax-n)(2t)V+2nxt=0.$
That is, we will establish that the equation
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}$
is equivalent to finding $y,z$ as solutions of
$V^{2}-(ax-n)(2t)V+2nxt=0.$
9.3 Preliminary Theorem
Before proving this equivalence, let us state the following theorem:
Theorem (Vieta’s Formulas – Equivalence Between Coefficients and Roots of a Quadratic Polynomial)
Let $X_{1}$ and $X_{2}$ be two real (or complex) numbers, and consider the quadratic equation:
$X^{2}-SX+P=0,$
where $S$ and $P$ are two real(or complex) parameters.
Then, the following two statements are equivalent:
1. $X_{1}$ and $X_{2}$ are solutions of the quadratic equation above.
2. $X_{1}$ and $X_{2}$ satisfy the relations:
$X_{1}+X_{2}=S,\quad X_{1}\cdot X_{2}=P.$
This property is a direct consequence of Vieta’s formulas.
9.4 Proof of the Equivalence
Let $a\in\mathbb{N}$ with $a\geq 4$, $n\in\mathbb{N}$ with $n\geq 2$, and $x,y,z\in\mathbb{N}^{*}$. We have:
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}$
which is equivalent to:
$\frac{a}{n}-\frac{1}{x}=\frac{1}{y}+\frac{1}{z}$
which is also equivalent to:
$\frac{ax-n}{nx}=\frac{y+z}{yz}.$
For $\frac{y+z}{yz}>0$, it is necessary that $ax-n>0$, i.e., $x\geq\left\lceil\frac{n}{a}\right\rceil+1$.
This last equation is equivalent to:
$yz(ax-n)=nx(y+z).$
Indeed, for any $a,c\in\mathbb{R}$ and $b,d\in\mathbb{R}^{*}$, we have:
$\frac{a}{b}=\frac{c}{d}\quad\Longleftrightarrow\quad ad=bc.$
(This is a fundamental property of fractions.)
Thus, our last equation is equivalent to the following system:
$(S)\quad\begin{cases}y+z=(ax-n)(2t),\\ yz=nx(2t),\end{cases}$
for any $t\in\mathbb{N}^{*}$.
By substituting $y+z$ with $(ax-n)(2t)$ and $yz$ with $nx(2t)$ in $yz(ax-n)=nx(y+z)$, the equation remains valid, which shows that this equation implies the system $(S)$.
9.5 Reciprocal Proof
Conversely, let us show that if:
$\begin{cases}y+z=(ax-n)(2t),\\ yz=nx(2t),\end{cases}$
for any $t\in\mathbb{N}^{*}$, then it follows that $yz(ax-n)=nx(y+z)$.
Indeed, since $y+z=(ax-n)(2t)$, we obtain:
$2t=\frac{y+z}{ax-n},$
noting that $ax-n\neq 0$ since we imposed $x\geq\left\lceil\frac{n}{a}\right\rceil+1$ to ensure $ax-n>0$.
On the other hand, since $yz=nx(2t)$, we also have:
$2t=\frac{yz}{nx}.$
By the transitivity of equality in $\mathbb{R}$, we deduce:
$\frac{y+z}{ax-n}=\frac{yz}{nx},$
which implies:
$nx(y+z)=yz(ax-n).$
This proves that system $(S)$ implies $nx(y+z)=yz(ax-n)$.
From the previous paragraph, we have justified the equivalence between $nx(y+z)=yz(ax-n)$ and the system $(S)$.
Finally, by applying Vieta’s theorem stated earlier, we conclude that the system $(S)$ is equivalent to stating that $y$ and $z$ are solutions of the following quadratic equation in $V$:
$V^{2}-(y+z)V+yz=0,$
which simplifies to:
$V^{2}-(ax-n)(2t)V+2nxt=0.$
This completes the proof of the equivalence.
10 Summary of the generalized Erdős-Straus and Sierpiński conjecture
We have shown that the generalized Erdős-Straus and Sierpiński conjecture (which includes the classical Erdős-Straus conjecture and the classical Sierpiński conjecture) is true if and only if the quadratic equation:
$V^{2}-(ax-n)(2t)V+2nxt=0$ (20)
always has solutions in $\mathbb{N}^{*}$ for any $a\in\mathbb{N}$ with $a\geq 4$ and any $n\in\mathbb{N}$ with $n\geq 2$. This reduces the difficulty of proving the conjecture to finding integer solutions (in $\mathbb{N}^{*}$) for a second-degree equation, whose discriminant is given by:
$\Delta=4t^{2}(ax-n)^{2}-8ntx.$ (21)
Rewriting,
$\Delta=4(t^{2}(ax-n)^{2}-2ntx).$ (22)
Thus, the proof of the generalized Erdős-Straus and Sierpiński conjecture can be reduced to verifying the following two conditions or refuting one of them to disprove the conjecture:
  1. 1.
    $t^{2}(ax-n)^{2}-2ntx$ is always a perfect square for all $(a,n)$.
  2. 2.
    $02t(ax-n)\pm\sqrt{\Delta}>$ 0. because condition (1) is sufficient, to obtain integer solutions, since $2t(ax-n)$ and $\sqrt{\Delta}$ will always be even, if (1) is verified, and the factor 2 in the formula will be simplified during the calculation of :
    $V=\frac{2t(ax-n)\pm\sqrt{\Delta}}{2}$ (23)
    is valid.
As in previous cases, we propose a partial resolution of the conjecture. We seek to solve the equation:
$V^{2}-(ax-n)(2t)V+2nxt=0,$ (24)
with discriminant:
$\Delta=4t^{2}(ax-n)^{2}-8ntx.$ (25)
Expanding,
$\Delta=(4a^{2}t^{2})x^{2}-(8ant^{2}+8nt)x+4t^{2}n^{2}.$ (26)
Solving $\Delta>0$, we obtain the solution range:
$x\in\left]0,\frac{n(at+1-\sqrt{2at+1})}{a^{2}t}\right[\cup\left]\frac{n(at+1+\sqrt{2at+1})}{a^{2}t},+\infty\right[\quad\text{with }x\in\mathbb{N}.$ (27)
similary by doing the same manipulations for the x for which $\Delta>0$
Since:
$\Delta=4t^{2}(ax-n)^{2}-8ntx=4t(t(ax-n)^{2}-2nx),$ (28)
we set:
$t=k=t(ax-n)^{2}-2nx,$ (29)
which leads to:
$k=\frac{2nx}{(ax-n)^{2}-1},$ (30)
where $k\in\mathbb{N}^{*}$ if $ax-n-1$ and $ax-n+1$ divide $2nx$.
Since:
$\Delta=4k^{2}=(2k)^{2},$ (31)
we obtain:
$V=\frac{(ax-n)(2k)\pm 2k}{2}.$ (32)
Thus,
$y=k(ax-n-1),\quad z=k(ax-n+1).$ (33)
Replacing $k$ by its expression:
$y=\frac{2nx}{ax-n+1},\quad z=\frac{2nx}{ax-n-1}.$ (34)
Verifying,
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$ (35)
As previously, the condition:
$x\geq\left\lceil\frac{n+1}{a}\right\rceil+1$ (36)
must hold. We will provide numerical computations at the end of this paper to generate certain $(a,n)$ pairs that satisfy this condition.
Similarly, for the second formula:
$\Delta=4(t^{2}(ax-n)^{2}-2ntx),$ (37)
for $\Delta>0$, we can set:
$t^{2}(ax-n)^{2}-2ntx=q^{2},\quad q\in\mathbb{N}.$
so that:
$\Delta=4q^{2}=(2q)^{2}.$ (38)
We will now prove that the conjecture is true under the condition that this perfect square exists. We obtain:
$y=\frac{2t(ax-n)-2q}{2}=t(ax-n)-q,\quad z=t(ax-n)+q.$ (39)
Verifying,
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$ (40)
Replacing $q^{2}$ by $t^{2}(ax-n)^{2}-2ntx$.
we will also provide numerical computations of some $(a,n)$ pairs satisfying this condition.
Summary:
we have demonstrated that for all $a\in\mathbb{N}$ with $a\geq 4$ and for all $n\in\mathbb{N}$ with $n\geq 2$, if there exists $x_{0}\in\mathbb{N}^{*}$ such that $ax_{0}-n-1$ and $ax_{0}-n+1$ divide $2nx_{0}$, then there exist $y$ and $z$ in $\mathbb{N}^{*}$ such that:
$y=\frac{2nx_{0}}{ax_{0}-n-1},\quad z=\frac{2nx_{0}}{ax_{0}-n+1}$
and we have:
$\frac{a}{n}=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z}.$
partially proving the conjecture under this divisibility condition.
Similarly, the second formula allows us to state that for any $a\in\mathbb{N}$ with $a\geq 4$ and for any $n\in\mathbb{N}$ with $n\geq 2$, if there exist $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that:
$t_{0}^{2}(ax_{0}-n)^{2}-2nt_{0}x_{0}=q^{2},\quad q\in\mathbb{N}.$
then there exist $y,z\in\mathbb{N}^{*}$ given by:
$y=t_{0}(ax_{0}-n)-q,\quad z=t_{0}(ax_{0}-n)+q$
such that:
$\frac{a}{n}=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z}.$
This also proves the conjecture under this condition.
I once again pose the same questions as for the classical Erdős-Straus conjecture and Sierpiński’s conjecture to the mathematical community in order to definitively resolve the generalized Erdős-Straus and Sierpiński’s conjecture.
11 Comparison with Mordell’s Approach
Mordell’s method provided a partial resolution of the Erdős-Straus conjecture by demonstrating that for all integers $n$ such that $n\not\equiv 1,11^{2},13^{2},17^{2},19^{2},23^{2}\pmod{840}$, the conjecture holds. However, for the exceptional cases left unresolved by Mordell, our second formula proves to be a powerful tool for generating solutions.
Through numerical testing, we observed that by naively programming the second formula, we can capture a significant portion of the exceptional cases left by Mordell. Specifically:
  • Among the values of $n$ in the range $[2,2000]$, we capture 86% of the cases not covered by Mordell.
11.1 Testing Second Formula for Mordell’s Exceptional Cases between [2, 2000]
illustration (Here, $L$ calculates the ratio $\frac{4}{n}$ and $R$ calculates the sum $\left(\frac{1}{x}+\frac{1}{y}+\frac{1}{z}\right)$, and the code outputs true if $R=L$ and false otherwise). Mordell’s exceptional cases in the given range : $121,169,289,361,529,841,961,1009,1129,1201,1369,1681,1801,1849,1969$ Total exceptional cases: 15
For $n=121$ (Mordell case), $x=33$, $y=726$, $z=726$, $t=66$: $L=0.0331$, $R=0.0331$, Conjecture holds: True
For $n=169$ (Mordell case), $x=44$, $y=1690$, $z=2860$, $t=325$: $L=0.0237$, $R=0.0237$, Conjecture holds: True
For $n=289$ (Mordell case), $x=85$, $y=578$, $z=2890$, $t=34$: $L=0.0138$, $R=0.0138$, Conjecture holds: True
For $n=361$ (Mordell case), $x=92$, $y=8664$, $z=10488$, $t=1368$: $L=0.0111$, $R=0.0111$, Conjecture holds: True
For $n=529$ (Mordell case), $x=138$, $y=6348$, $z=6348$, $t=276$: $L=0.0076$, $R=0.0076$, Conjecture holds: True
For $n=841$ (Mordell case), $x=220$, $y=6380$, $z=18502$, $t=319$: $L=0.0048$, $R=0.0048$, Conjecture holds: True
For $n=961$ (Mordell case), $x=248$, $y=15376$, $z=15376$, $t=496$: $L=0.0042$, $R=0.0042$, Conjecture holds: True
For $n=1009$ (Mordell case), $x=5676$, $y=264$, $z=3818056$, $t=88$: $L=0.0040$, $R=0.0040$, Conjecture holds: True
For $n=1129$ (Mordell case), $x=3952$, $y=304$, $z=2230904$, $t=76$: $L=0.0035$, $R=0.0035$, Conjecture holds: True
For $n=1369$ (Mordell case), $x=348$, $y=38332$, $z=45066$, $t=1813$: $L=0.0029$, $R=0.0029$, Conjecture holds: True
For $n=1681$ (Mordell case), $x=423$, $y=121032$, $z=138744$, $t=11808$: $L=0.0024$, $R=0.0024$, Conjecture holds: True
For $n=1849$ (Mordell case), $x=473$, $y=40678$, $z=40678$, $t=946$: $L=0.0022$, $R=0.0022$, Conjecture holds: True
For $n=1969$ (Mordell case), $x=495$, $y=177210$, $z=177210$, $t=16110$: $L=0.0020$, $R=0.0020$, Conjecture holds: True
Percentage of Mordell’s exceptional cases verified: 86.67%
  • In the range $[2000,5000]$, we capture 60% of the remaining cases.
11.2 Testing Second Formula for Mordell’s Exceptional Cases: between [2000, 5000]
Mordell’s exceptional cases in the given range:
$\begin{array}[]{c}2041,2209,2521,2641,2689,2809,2881,3049,3361,3481,\\ 3529,3649,3721,3889,4201,4321,4369,4489,4561,4729\\ \end{array}$
Total exceptional cases: 20
Testing Second Formula for Mordell’s Exceptional Cases:
$\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr n&x&y&z&t&L&R&Conjectureholds\\ \hline\cr 2041&520&48984&61230&1413&0.0020&0.0020&True\\ 2209&564&53016&53016&1128&0.0018&0.0018&True\\ 2641&665&184870&184870&9730&0.0015&0.0015&True\\ 2809&710&112360&150520&4240&0.0014&0.0014&True\\ 2881&731&97954&97954&2278&0.0014&0.0014&True\\ 3481&885&104430&104430&1770&0.0011&0.0011&True\\ 3649&984&21894&29192&89&0.0011&0.0011&True\\ 3721&976&29768&59536&244&0.0011&0.0011&True\\ 4201&1056&369688&403296&16804&0.0010&0.0010&True\\ 4321&1085&483952&503440&25984&0.0009&0.0009&True\\ 4369&1104&200974&209712&4369&0.0009&0.0009&True\\ 4489&1139&152626&152626&2278&0.0009&0.0009&True\\ \hline\cr\end{array}$
Percentage of Mordell’s exceptional cases verified: 60.00%
In the range $[5000,10000]$, we capture 61% of the remaining cases.
11.3 Testing Second Formula for Mordell’s Exceptional Cases : between [5000, 10000]
Mordell’s exceptional cases in the given range:
$\begin{array}[]{c}5041,5161,5209,5329,5401,5569,5881,6001,6049,6169,\\ 6241,6409,6721,6841,6889,7009,7081,7249,7561,7681,\\ 7729,7849,7921,8089,8401,8521,8569,8689,8761,8929,\\ 9241,9361,9409,9529,9601,9769\end{array}$
Total exceptional cases : 36
Testing Second Formula for Mordell’s Exceptional Cases:
$\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr n&x&y&z&t&L&R&Conjectureholds\\ \hline\cr 5041&1278&181476&181476&2556&0.0008&0.0008&True\\ 5161&1314&115924&185796&1588&0.0008&0.0008&True\\ 5329&1460&21316&53290&73&0.0008&0.0008&True\\ 5401&1353&1328646&1328646&120786&0.0007&0.0007&True\\ 6001&1765&12002&60010&34&0.0007&0.0007&True\\ 6049&1518&798468&798468&34716&0.0007&0.0007&True\\ 6169&1544&2680384&2763712&388864&0.0006&0.0006&True\\ 6241&1580&249640&249640&3160&0.0006&0.0006&True\\ 6409&1612&512720&548080&13600&0.0006&0.0006&True\\ 6721&1683&2056626&2056626&186966&0.0006&0.0006&True\\ 6889&1743&289338&289338&3486&0.0006&0.0006&True\\ 7009&1763&574738&574738&13366&0.0006&0.0006&True\\ 7081&1848&74690&96360&275&0.0006&0.0006&True\\ 7249&1815&2392170&2392170&217470&0.0006&0.0006&True\\ 7729&1947&510114&510114&8646&0.0005&0.0005&True\\ 7849&1974&659316&659316&14028&0.0005&0.0005&True\\ 7921&2136&47526&63368&89&0.0005&0.0005&True\\ 8401&2108&1142536&1142536&36856&0.0005&0.0005&True\\ 8569&2145&3341910&3341910&303810&0.0005&0.0005&True\\ 9361&2343&3987786&3987786&362526&0.0004&0.0004&True\\ 9409&2716&18818&263452&97&0.0004&0.0004&True\\ 9529&2932&19058&38116&13&0.0004&0.0004&True\\ \hline\cr\\ \end{array}$
Percentage of Mordell’s exceptional cases verified: 61.11%
These results show that the second formula is a machine for generating solutions, including for the cases left open by Mordell(These results show that this formula naively captures 69% of the cases of $n$ in the range $[2,10000]$ that satisfy the conjecture but are not captured by Mordell’s formula). Moreover, even with a relatively simple computational implementation, a large number of solutions have been found, suggesting that a complete proof of the conjecture could be within reach.
Finally, it is worth noting that these statistics can be further refined by increasing the precision of the search for optimal parameters $x$ and $t$. This work thus opens up a new and promising perspective in the search for a complete proof of the Erdős-Straus conjecture.
To allow further verification and exploration beyond the values presented in this article, we include the implementation of our computational method in the appendix. This will enable researchers to test additional values of $n$ without overloading the main content of this paper.
In the following tables, we provide the numerical values of some $n$, $x_{0}$, $y$, and $z$ that satisfy the conditions of each formula. The value $L$ (for "Left") represents the ratio $a/n$, while $R$ (for "Right") is computed as $R=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z}$.
The last column verifies whether $L=R$:
  • Rows labeled "True" indicate equality, meaning the conjecture holds for the given values of $n$, $x_{0}$, $y$, and $z$.
  • Rows labeled "False" indicate that the conjecture does not hold for those specific values.
If a value of $n$ is not displayed in the table, this does not mean that the number does not satisfy the formula or the conjecture. It simply indicates that the algorithm did not find an $x_{0}$ and/or a $t_{0}$ within the specific search range that meets one of the formula’s conditions for that $n$. This is due to the fact that there are infinitely many possible choices for $x_{0}$.
To avoid overloading the article, I have included in the appendix the Python code for each formula so that interested readers can verify for themselves how these formulas generate solutions, especially the second formula, which captures almost all $n$ in a given interval (for $a=4$, the classical Erdős-Straus conjecture), and potentially all $n$, even for very large values, with a significant variety of solutions for each $n$. This can be seen in Tables 2,3,4,5 below, where we capture (for $a=4$) all $n$ in the range $[2,200]$ that satisfy the conjecture using the second formula. I strongly encourage readers to copy and paste the second code from the appendix, run it in a Python environment, and observe how the second formula generates huge solutions (capturing all $n$ for $a=4$ within an interval up to a very high threshold).You will notice when running this code that sometimes a single $n$ can generate around 40 distinct solutions $(x,y,z)$.
I had fun testing naively, numerically (using the fourth code in the appendix of this article) for:
$n\in[2,100],\quad n\in]100,200],\quad n\in]200,300],~{}\cdots,~{}n\in]1900,2000]$
searching for values of $n$ for which the polynomial :
$t^{2}(4x-n)^{2}-2nxt$
admits a perfect square while also identifying the non-captured $n$ (which I call the recalcitrant cases) and computing the percentage of captured $n$ per interval.
I obtained the following results:
For $n\in[2,2000]$ (excluding the subintervals $[500,600],[1200,1300],[1800,1900]$), no recalcitrant cases were found, achieving a 100% capture rate, meaning this polynomial admits a perfect square for all $n$ in these ranges.
For $n\in[500,600]$, a 99.01% capture rate was recorded, with a single recalcitrant case at $n=577$. I handled this manually using the same reasoning proposed at the beginning of this article: Since
$x\geq\left\lceil\frac{577}{4}\right\rceil+1=145,$
I took $x=145$, leading to the discriminant
$\Delta=t(9t-723002).$
Please refer to the fifth code in the appendix of this article to execute this case. You will find that for $t\in[25000,40000]$, we obtain $t=33466$, $\Delta=4479892624$, and $\sqrt{\Delta}=66932$.
For $n\in[1200,1300]$, a 97.03% capture rate was recorded, with three recalcitrant cases:
$n\in\{1201,1249,1277\}.$
For $n\in[1800,1900]$, a 99.01% capture rate was recorded, with a single recalcitrant case at $n=1801$.
For the cases $n\in\{1249,1277\}$, one can proceed in the same way as for $n=577$.
However, the cases $n\in\{1201,1801\}$ are exactly the ones not proven by Mordell and were also not captured in our previous analysis over the range $[2,2000]$, where 15 cases were analyzed, 13 were captured (86%), and the two non-captured cases were also $1201$ and $1801$.
It seems that these two cases require a more refined analysis, which I leave to the mathematical community, as I did not have enough time to focus on them.
(You may also attempt to handle them similarly to $n=577$ by using the fifth code in the appendix to extend the search range for $t$ and $x$. This is just a suggested approach, but I cannot guarantee its effectiveness, as we have infinitely many choices for $t$ and $x$. Therefore, a more detailed analytical study is probably required for these two cases).
Similarly, I had fun using the same code to naively verify whether the polynomial
$t^{2}(ax-n)^{2}-2nxt$
always admits a perfect square for any fixed $a\geq 5$ and for all $n\geq 2$.
What I observed is that, initially, there are some recalcitrant cases. However, these recalcitrant cases tend to decrease as $n$ increases.
In summary, naively, we observe approximately 95% of captured cases.
However, to reduce the number of recalcitrant cases, the reader may proceed as in the case of $n=577$ previously discussed for $a=4$.
Based on the preceding observations, we numerically observe that the property appears to hold for a very large number of values of $n$. If this regularity is confirmed, one could then formulate it as a conjecture, which we propose here as an open question for the mathematical community.
Conjecture 1: For all $n\in\mathbb{N}$ with $n\geq 2$, there exist $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that
$t_{0}^{2}(4x_{0}-n)^{2}-2nx_{0}t_{0}=q^{2},\quad q\in\mathbb{N}.$
Conjecture 2: For all $a\in\mathbb{N}$ with $a\geq 5$, and for all $n\in\mathbb{N}$ with $n\geq 2$, there exists $N_{0}\in\mathbb{N}$ such that for all $n\geq N_{0}$, there exist $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that
$t_{0}^{2}(ax_{0}-n)^{2}-2nx_{0}t_{0}=q^{2},\quad q\in\mathbb{N}.$
Table 1: Results with the first formula of the Erdős-Straus conjecture test for n from 2 to 90
n x y z L R L = R (verification)
3 2 3 2 1.3333 1.3333 True
5 2 10 5 0.8000 0.8000 True
6 2 24 8 0.6667 0.6667 True
9 4 12 9 0.4444 0.4444 True
10 3 60 20 0.4000 0.4000 True
13 4 52 26 0.3077 0.3077 True
14 5 28 20 0.2857 0.2857 True
18 5 180 60 0.2222 0.2222 True
19 6 57 38 0.2105 0.2105 True
20 6 80 48 0.2000 0.2000 True
21 6 126 63 0.1905 0.1905 True
22 6 264 88 0.1818 0.1818 True
27 8 108 72 0.1481 0.1481 True
28 9 72 56 0.1429 0.1429 True
29 8 232 116 0.1379 0.1379 True
30 8 480 160 0.1333 0.1333 True
34 9 612 204 0.1176 0.1176 True
35 12 70 60 0.1143 0.1143 True
36 10 240 144 0.1111 0.1111 True
37 10 370 185 0.1081 0.1081 True
39 16 52 48 0.1026 0.1026 True
41 12 164 123 0.0976 0.0976 True
42 11 924 308 0.0952 0.0952 True
43 12 258 172 0.0930 0.0930 True
45 12 540 270 0.0889 0.0889 True
46 12 1104 368 0.0870 0.0870 True
49 16 112 98 0.0816 0.0816 True
50 14 280 200 0.0800 0.0800 True
51 14 357 238 0.0784 0.0784 True
53 14 742 371 0.0755 0.0755 True
54 14 1512 504 0.0741 0.0741 True
55 16 220 176 0.0727 0.0727 True
56 15 560 336 0.0714 0.0714 True
57 16 304 228 0.0702 0.0702 True
58 15 1740 580 0.0690 0.0690 True
60 16 640 384 0.0667 0.0667 True
61 16 976 488 0.0656 0.0656 True
66 17 2244 748 0.0606 0.0606 True
67 18 603 402 0.0597 0.0597 True
69 18 1242 621 0.0580 0.0580 True
70 18 2520 840 0.0571 0.0571 True
71 20 355 284 0.0563 0.0563 True
75 20 750 500 0.0533 0.0533 True
77 20 1540 770 0.0519 0.0519 True
78 20 3120 1040 0.0513 0.0513 True
80 21 1120 672 0.0500 0.0500 True
82 21 3444 1148 0.0488 0.0488 True
85 22 1870 935 0.0471 0.0471 True
89 24 712 534 0.0449 0.0449 True
90 23 4140 1380 0.0444 0.0444 True
Table 2: Results of Erdős–Straus Conjecture with the second formula testing for n from 2 to 50
n x y z L R L = R (verification)
2 1 2 2 2.0 2.0 True
3 1 6 6 1.3333 1.3333 True
4 2 4 4 1.0 1.0 True
5 2 4 20 0.8 0.8 True
6 2 12 12 0.6667 0.6667 True
7 2 28 28 0.5714 0.5714 True
8 3 12 12 0.5 0.5 True
9 3 18 18 0.4444 0.4444 True
10 3 30 30 0.4 0.4 True
11 3 66 66 0.3636 0.3636 True
12 4 24 24 0.3333 0.3333 True
13 4 26 52 0.3077 0.3077 True
14 4 56 56 0.2857 0.2857 True
15 4 120 120 0.2667 0.2667 True
16 5 40 40 0.25 0.25 True
17 5 34 170 0.2353 0.2353 True
18 5 90 90 0.2222 0.2222 True
19 5 190 190 0.2105 0.2105 True
20 6 60 60 0.2 0.2 True
21 6 84 84 0.1905 0.1905 True
22 6 132 132 0.1818 0.1818 True
23 6 276 276 0.1739 0.1739 True
24 7 84 84 0.1667 0.1667 True
25 7 100 140 0.16 0.16 True
26 7 182 182 0.1538 0.1538 True
27 7 378 378 0.1481 0.1481 True
28 8 112 112 0.1429 0.1429 True
29 8 116 232 0.1379 0.1379 True
30 8 240 240 0.1333 0.1333 True
31 8 496 496 0.129 0.129 True
32 9 144 144 0.125 0.125 True
33 9 198 198 0.1212 0.1212 True
34 9 306 306 0.1176 0.1176 True
35 9 630 630 0.1143 0.1143 True
36 10 180 180 0.1111 0.1111 True
37 10 148 740 0.1081 0.1081 True
38 10 380 380 0.1053 0.1053 True
39 10 780 780 0.1026 0.1026 True
40 11 220 220 0.1 0.1 True
41 12 82 492 0.0976 0.0976 True
42 11 462 462 0.0952 0.0952 True
43 11 946 946 0.093 0.093 True
44 12 264 264 0.0909 0.0909 True
45 12 360 360 0.0889 0.0889 True
46 12 552 552 0.087 0.087 True
47 12 1128 1128 0.0851 0.0851 True
48 13 312 312 0.0833 0.0833 True
49 14 196 196 0.0816 0.0816 True
50 13 650 650 0.08 0.08 True
Table 3: Results of Erdős–Straus Conjecture with the second formula testing for n from 51 to 100
n x y z L R L = R (verification)
51 13 1326 1326 0.0784 0.0784 True
52 14 364 364 0.0769 0.0769 True
53 16 80 2120 0.0755 0.0755 True
54 14 756 756 0.0741 0.0741 True
55 14 1540 1540 0.0727 0.0727 True
56 15 420 420 0.0714 0.0714 True
57 15 570 570 0.0702 0.0702 True
58 15 870 870 0.069 0.069 True
59 15 1770 1770 0.0678 0.0678 True
60 16 480 480 0.0667 0.0667 True
61 16 488 976 0.0656 0.0656 True
62 16 992 992 0.0645 0.0645 True
63 16 2016 2016 0.0635 0.0635 True
64 17 544 544 0.0625 0.0625 True
65 17 650 850 0.0615 0.0615 True
66 17 1122 1122 0.0606 0.0606 True
67 17 2278 2278 0.0597 0.0597 True
68 18 612 612 0.0588 0.0588 True
69 18 828 828 0.058 0.058 True
70 18 1260 1260 0.0571 0.0571 True
71 18 2556 2556 0.0563 0.0563 True
72 19 684 684 0.0556 0.0556 True
73 20 292 730 0.0548 0.0548 True
74 19 1406 1406 0.0541 0.0541 True
75 19 2850 2850 0.0533 0.0533 True
76 20 760 760 0.0526 0.0526 True
77 20 980 1078 0.0519 0.0519 True
78 20 1560 1560 0.0513 0.0513 True
79 20 3160 3160 0.0506 0.0506 True
80 21 840 840 0.05 0.05 True
81 21 1134 1134 0.0494 0.0494 True
82 21 1722 1722 0.0488 0.0488 True
83 21 3486 3486 0.0482 0.0482 True
84 22 924 924 0.0476 0.0476 True
85 23 340 1564 0.0471 0.0471 True
86 22 1892 1892 0.0465 0.0465 True
87 22 3828 3828 0.046 0.046 True
88 23 1012 1012 0.0455 0.0455 True
89 24 534 712 0.0449 0.0449 True
90 23 2070 2070 0.0444 0.0444 True
91 23 4186 4186 0.044 0.044 True
92 24 1104 1104 0.0435 0.0435 True
93 24 1488 1488 0.043 0.043 True
94 24 2256 2256 0.0426 0.0426 True
95 24 4560 4560 0.0421 0.0421 True
96 25 1200 1200 0.0417 0.0417 True
97 28 194 2716 0.0412 0.0412 True
98 25 2450 2450 0.0408 0.0408 True
99 25 4950 4950 0.0404 0.0404 True
100 26 1300 1300 0.04 0.04 True
Table 4: Results of Erdős–Straus Conjecture with second formula testing for $n$ from 101 to 150
n x y z L R L = R (verification)
101 38 76 7676 0.0396 0.0396 True
102 26 2652 2652 0.0392 0.0392 True
103 26 5356 5356 0.0388 0.0388 True
104 27 1404 1404 0.0385 0.0385 True
105 27 1890 1890 0.0381 0.0381 True
106 27 2862 2862 0.0377 0.0377 True
107 27 5778 5778 0.0374 0.0374 True
108 28 1512 1512 0.037 0.037 True
109 28 1526 3052 0.0367 0.0367 True
110 28 3080 3080 0.0364 0.0364 True
111 28 6216 6216 0.036 0.036 True
112 29 1624 1624 0.0357 0.0357 True
113 132 36 22374 0.0354 0.0354 True
114 29 3306 3306 0.0351 0.0351 True
115 29 6670 6670 0.0348 0.0348 True
116 30 1740 1740 0.0345 0.0345 True
117 30 2340 2340 0.0342 0.0342 True
118 30 3540 3540 0.0339 0.0339 True
119 30 7140 7140 0.0336 0.0336 True
120 31 1860 1860 0.0333 0.0333 True
121 33 726 726 0.0331 0.0331 True
122 31 3782 3782 0.0328 0.0328 True
123 31 7626 7626 0.0325 0.0325 True
124 32 1984 1984 0.0323 0.0323 True
125 32 2400 3000 0.032 0.032 True
126 32 4032 4032 0.0317 0.0317 True
127 32 8128 8128 0.0315 0.0315 True
128 33 2112 2112 0.0312 0.0312 True
129 33 2838 2838 0.031 0.031 True
130 33 4290 4290 0.0308 0.0308 True
131 33 8646 8646 0.0305 0.0305 True
132 34 2244 2244 0.0303 0.0303 True
133 34 2856 3192 0.0301 0.0301 True
134 34 4556 4556 0.0299 0.0299 True
135 34 9180 9180 0.0296 0.0296 True
136 35 2380 2380 0.0294 0.0294 True
137 35 2740 3836 0.0292 0.0292 True
138 35 4830 4830 0.029 0.029 True
139 35 9730 9730 0.0288 0.0288 True
140 36 2520 2520 0.0286 0.0286 True
141 36 3384 3384 0.0284 0.0284 True
142 36 5112 5112 0.0282 0.0282 True
143 36 10296 10296 0.028 0.028 True
144 37 2664 2664 0.0278 0.0278 True
145 37 3190 4070 0.0276 0.0276 True
146 37 5402 5402 0.0274 0.0274 True
147 37 10878 10878 0.0272 0.0272 True
148 38 2812 2812 0.027 0.027 True
149 70 80 16688 0.0268 0.0268 True
150 38 5700 5700 0.0267 0.0267 True
Table 5: Results of Erdős–Straus Conjecture with second formula testing for n from 150 to 200
n x y z L R L = R (verification)
151 38 11476 11476 0.0265 0.0265 True
152 39 2964 2964 0.0263 0.0263 True
153 39 3978 3978 0.0261 0.0261 True
154 39 6006 6006 0.026 0.026 True
155 39 12090 12090 0.0258 0.0258 True
156 40 3120 3120 0.0256 0.0256 True
157 40 3768 4710 0.0255 0.0255 True
158 40 6320 6320 0.0253 0.0253 True
159 40 12720 12720 0.0252 0.0252 True
160 41 3280 3280 0.025 0.025 True
161 42 1932 1932 0.0248 0.0248 True
162 41 6642 6642 0.0247 0.0247 True
163 41 13366 13366 0.0245 0.0245 True
164 42 3444 3444 0.0244 0.0244 True
165 42 4620 4620 0.0242 0.0242 True
166 42 6972 6972 0.0241 0.0241 True
167 42 14028 14028 0.024 0.024 True
168 43 3612 3612 0.0238 0.0238 True
169 44 1690 2860 0.0237 0.0237 True
170 43 7310 7310 0.0235 0.0235 True
171 43 14706 14706 0.0234 0.0234 True
172 44 3784 3784 0.0233 0.0233 True
173 45 1730 3114 0.0231 0.0231 True
174 44 7656 7656 0.023 0.023 True
175 44 15400 15400 0.0229 0.0229 True
176 45 3960 3960 0.0227 0.0227 True
177 45 5310 5310 0.0226 0.0226 True
178 45 8010 8010 0.0225 0.0225 True
179 45 16110 16110 0.0223 0.0223 True
180 46 4140 4140 0.0222 0.0222 True
181 48 1086 2896 0.0221 0.0221 True
182 46 8372 8372 0.022 0.022 True
183 46 16836 16836 0.0219 0.0219 True
184 47 4324 4324 0.0217 0.0217 True
185 47 5180 6580 0.0216 0.0216 True
186 47 8742 8742 0.0215 0.0215 True
187 47 17578 17578 0.0214 0.0214 True
188 48 4512 4512 0.0213 0.0213 True
189 48 6048 6048 0.0212 0.0212 True
190 48 9120 9120 0.0211 0.0211 True
191 48 18336 18336 0.0209 0.0209 True
192 49 4704 4704 0.0208 0.0208 True
193 676 52 65234 0.0207 0.0207 True
194 49 9506 9506 0.0206 0.0206 True
195 49 19110 19110 0.0205 0.0205 True
196 50 4900 4900 0.0204 0.0204 True
197 148 74 29156 0.0203 0.0203 True
198 50 9900 9900 0.0202 0.0202 True
199 50 19900 19900 0.0201 0.0201 True
200 51 5100 5100 0.02 0.02 True
Table 6: Results of Sierpiński Conjecture with second formula testing for $n$ from $2$ to $50$
n x y z L R L = R (verification)
3 1 2 6 1.6667 1.6667 True
4 1 8 8 1.25 1.25 True
6 2 4 12 0.8333 0.8333 True
7 5 2 70 0.7143 0.7143 True
8 2 16 16 0.625 0.625 True
9 2 36 36 0.5556 0.5556 True
11 5 4 220 0.4545 0.4545 True
12 3 24 24 0.4167 0.4167 True
13 3 26 78 0.3846 0.3846 True
14 3 84 84 0.3571 0.3571 True
15 4 20 30 0.3333 0.3333 True
16 4 32 32 0.3125 0.3125 True
17 4 34 68 0.2941 0.2941 True
18 4 72 72 0.2778 0.2778 True
19 4 152 152 0.2632 0.2632 True
20 5 40 40 0.25 0.25 True
21 5 42 70 0.2381 0.2381 True
22 5 44 220 0.2273 0.2273 True
23 5 70 322 0.2174 0.2174 True
24 5 240 240 0.2083 0.2083 True
25 6 60 60 0.2 0.2 True
26 6 52 156 0.1923 0.1923 True
27 6 108 108 0.1852 0.1852 True
28 6 168 168 0.1786 0.1786 True
29 6 348 348 0.1724 0.1724 True
30 7 60 140 0.1667 0.1667 True
31 7 62 434 0.1613 0.1613 True
32 7 140 160 0.1562 0.1562 True
33 7 176 336 0.1515 0.1515 True
34 7 476 476 0.1471 0.1471 True
35 8 70 280 0.1429 0.1429 True
36 8 144 144 0.1389 0.1389 True
37 8 148 296 0.1351 0.1351 True
38 8 304 304 0.1316 0.1316 True
39 8 624 624 0.1282 0.1282 True
40 9 120 180 0.125 0.125 True
41 26 12 6396 0.122 0.122 True
42 9 252 252 0.119 0.119 True
43 9 258 774 0.1163 0.1163 True
44 9 792 792 0.1136 0.1136 True
45 10 180 180 0.1111 0.1111 True
46 10 140 644 0.1087 0.1087 True
47 10 188 940 0.1064 0.1064 True
48 10 480 480 0.1042 0.1042 True
49 10 980 980 0.102 0.102 True
50 11 220 220 0.1 0.1 True
Table 7: Results of Sierpiński Conjecture with second formula testing for $n$ from $51$ to $81$
n x y z L R L = R (verification)
51 12 102 204 0.098 0.098 True
52 11 352 416 0.0962 0.0962 True
53 12 106 636 0.0943 0.0943 True
54 11 1188 1188 0.0926 0.0926 True
55 12 220 330 0.0909 0.0909 True
56 12 336 336 0.0893 0.0893 True
57 12 456 456 0.0877 0.0877 True
58 12 696 696 0.0862 0.0862 True
59 12 1416 1416 0.0847 0.0847 True
60 13 260 390 0.0833 0.0833 True
61 95 14 81130 0.082 0.082 True
62 14 124 868 0.0806 0.0806 True
63 13 728 936 0.0794 0.0794 True
64 13 1664 1664 0.0781 0.0781 True
65 14 280 520 0.0769 0.0769 True
66 14 352 672 0.0758 0.0758 True
67 313 14 293594 0.0746 0.0746 True
68 14 952 952 0.0735 0.0735 True
69 14 1932 1932 0.0725 0.0725 True
70 15 420 420 0.0714 0.0714 True
71 15 426 710 0.0704 0.0704 True
72 15 720 720 0.0694 0.0694 True
73 15 876 1460 0.0685 0.0685 True
74 15 2220 2220 0.0676 0.0676 True
75 16 480 480 0.0667 0.0667 True
76 16 608 608 0.0658 0.0658 True
77 16 770 880 0.0649 0.0649 True
78 16 1248 1248 0.0641 0.0641 True
79 16 2528 2528 0.0633 0.0633 True
80 17 400 850 0.0625 0.0625 True
81 18 324 324 0.0617 0.0617 True
Table 8: Results of the generalized conjecture ($a=8$ )with second formula Testing for $n$ from $2$ to $50$
n x y z L R L = R (verification)
5 1 2 10 1.6 1.6 True
6 1 6 6 1.3333 1.3333 True
7 1 14 14 1.1429 1.1429 True
10 2 4 20 0.8 0.8 True
12 2 12 12 0.6667 0.6667 True
13 9 2 234 0.6154 0.6154 True
14 2 28 28 0.5714 0.5714 True
15 2 60 60 0.5333 0.5333 True
16 3 8 24 0.5 0.5 True
18 3 18 18 0.4444 0.4444 True
19 3 12 228 0.4211 0.4211 True
20 3 24 40 0.4 0.4 True
21 3 42 42 0.381 0.381 True
22 3 66 66 0.3636 0.3636 True
23 3 138 138 0.3478 0.3478 True
24 4 24 24 0.3333 0.3333 True
25 4 20 50 0.32 0.32 True
26 5 10 130 0.3077 0.3077 True
27 4 36 54 0.2963 0.2963 True
28 4 56 56 0.2857 0.2857 True
29 4 58 116 0.2759 0.2759 True
30 4 120 120 0.2667 0.2667 True
31 4 248 248 0.2581 0.2581 True
32 5 40 40 0.25 0.25 True
33 5 30 110 0.2424 0.2424 True
34 5 34 170 0.2353 0.2353 True
35 5 70 70 0.2286 0.2286 True
36 5 72 120 0.2222 0.2222 True
37 5 74 370 0.2162 0.2162 True
38 5 190 190 0.2105 0.2105 True
39 5 390 390 0.2051 0.2051 True
40 6 48 80 0.2 0.2 True
41 6 36 1476 0.1951 0.1951 True
42 6 84 84 0.1905 0.1905 True
43 52 6 6708 0.186 0.186 True
44 6 132 132 0.1818 0.1818 True
45 6 180 180 0.1778 0.1778 True
46 6 276 276 0.1739 0.1739 True
47 6 564 564 0.1702 0.1702 True
48 7 56 168 0.1667 0.1667 True
49 7 98 98 0.1633 0.1633 True
50 7 100 140 0.16 0.16 True
Table 9: Results of the generalized conjecture ($a=8$ )with second formula Testing for $n$ from $51$ to $72$
n x y z L R L = R (verification)
51 7 102 238 0.1569 0.1569 True
52 7 140 260 0.1538 0.1538 True
53 8 40 1060 0.1509 0.1509 True
54 7 378 378 0.1481 0.1481 True
55 7 770 770 0.1455 0.1455 True
56 8 112 112 0.1429 0.1429 True
57 8 114 152 0.1404 0.1404 True
58 8 116 232 0.1379 0.1379 True
59 8 118 472 0.1356 0.1356 True
60 8 240 240 0.1333 0.1333 True
61 8 244 488 0.1311 0.1311 True
62 8 496 496 0.129 0.129 True
63 8 1008 1008 0.127 0.127 True
64 9 144 144 0.125 0.125 True
65 9 156 180 0.1231 0.1231 True
66 9 198 198 0.1212 0.1212 True
67 9 134 1206 0.1194 0.1194 True
68 9 204 612 0.1176 0.1176 True
69 9 414 414 0.1159 0.1159 True
70 9 630 630 0.1143 0.1143 True
71 9 1278 1278 0.1127 0.1127 True
72 10 144 240 0.1111 0.1111 True
Table 10: Results of the generalized conjecture ($a=11$) with first formula Testing for $n$ from $500$ to $625$
n x y z L R L = R (verification)
501 46 11523 7682 0.0219560878 0.0219560878 True
503 46 23138 11569 0.0218687873 0.0218687873 True
504 46 46368 15456 0.0218253968 0.0218253968 True
510 49 1785 1666 0.0215686275 0.0215686275 True
513 50 1425 1350 0.0214424951 0.0214424951 True
514 47 24158 12079 0.0214007782 0.0214007782 True
521 48 8336 6252 0.0211132438 0.0211132438 True
523 48 12552 8368 0.0210325048 0.0210325048 True
525 48 25200 12600 0.0209523809 0.0209523809 True
526 48 50496 16832 0.0209125475 0.0209125475 True
528 51 1683 1584 0.0208333333 0.0208333333 True
531 49 7434 5782 0.0207156309 0.0207156309 True
534 49 13083 8722 0.0205992509 0.0205992509 True
536 49 26264 13132 0.0205223881 0.0205223881 True
537 49 52626 17542 0.0204841713 0.0204841713 True
545 51 3706 3270 0.0201834862 0.0201834862 True
546 50 18200 10920 0.0201465201 0.0201465201 True
547 50 27350 13675 0.0201096892 0.0201096892 True
550 51 5610 4675 0.0200000000 0.0200000000 True
556 51 14178 9452 0.0197841727 0.0197841727 True
558 51 28458 14229 0.0197132616 0.0197132616 True
559 51 57018 19006 0.0196779964 0.0196779964 True
567 52 14742 9828 0.0194003527 0.0194003527 True
569 52 29588 14794 0.0193321617 0.0193321617 True
570 52 59280 19760 0.0192982456 0.0192982456 True
575 54 3450 3105 0.0191304348 0.0191304348 True
576 53 10176 7632 0.0190972222 0.0190972222 True
580 53 30740 15370 0.0189655172 0.0189655172 True
581 54 5229 4482 0.0189328743 0.0189328743 True
584 55 3212 2920 0.0188356164 0.0188356164 True
589 54 15903 10602 0.0186757216 0.0186757216 True
590 54 21240 12744 0.0186440678 0.0186440678 True
591 54 31914 15957 0.0186125212 0.0186125212 True
592 54 63936 21312 0.0185810811 0.0185810811 True
594 55 6534 5445 0.0185185185 0.0185185185 True
596 55 8195 6556 0.0184563758 0.0184563758 True
600 55 16500 11000 0.0183333333 0.0183333333 True
601 56 4808 4207 0.0183028286 0.0183028286 True
602 55 33110 16555 0.0182724252 0.0182724252 True
603 55 66330 22110 0.0182421227 0.0182421227 True
609 56 11368 8526 0.0180623974 0.0180623974 True
610 56 13664 9760 0.0180327869 0.0180327869 True
612 65 780 765 0.0179738562 0.0179738562 True
613 56 34328 17164 0.0179445351 0.0179445351 True
620 57 11780 8835 0.0177419355 0.0177419355 True
622 57 17727 11818 0.0176848875 0.0176848875 True
624 57 35568 17784 0.0176282051 0.0176282051 True
Appendix: Python Code for Testing the Generalized Erdős-Straus Conjecture (First Formula)
The following code allows testing the generalized Erdős-Straus conjecture numerically for different values of $a$ and $n$. It provides direct output of the results without generating a CSV file.
[…]
2
3 def generalized_erdos_straus_conjecture(a, n_min, n_max):
4 """
[…]
6 This function prints results directly instead of saving them to a file.
7
8 :param a: Integer, the numerator in the equation a/n = 1/x + 1/y + 1/z.
9 :param n_min: Integer, the starting value of n.
10 :param n_max: Integer, the ending value of n.
11 """
12 epsilon = 1e-9 # Tolerance for floating-point comparison
13 for n in range(n_min, n_max + 1):
14 solution_found = False
15 L = a / n # Left-hand side of the equation
16 for x in range(n // a + 1, 10 * n): # Searching for x > n/a
17 lambda_value = a * x - n - 1
18 if lambda_value <= 0:
19 continue
20 y = (2 * n * x) // lambda_value
21 z = (2 * n * x) // (lambda_value + 2)
22
23 # Ensure y and z are integers
24 if (2 * n * x) % lambda_value == 0 and (2 * n * x) % (lambda_value + 2) == 0:
25 R = 1/x + 1/y + 1/z # Right-hand side of the equation
26 is_true = abs(L - R) < epsilon # Floating-point comparison with tolerance
27 print(f"For n = {n}, x = {x}, y = {y}, z = {z}: L = {L:.4f}, R = {R:.4f}, Conjecture holds: {is_true}")
28 solution_found = True
29 break
30
31 # Example usage:
32 a = 4 #(for the classic Erdos_Strauss Conjecture)
33 n_min = 2
34 n_max = 100
35 generalized_erdos_straus_conjecture(a, n_min, n_max)
Appendix: Python Code for Testing the Generalized Erdős-Straus Conjecture (Second Formula)
The following code tests the generalized Erdős-Straus conjecture using the second formula independently. It prints results directly based on the existence of a perfect square.
1 def second_formula_erdos_straus(a, n_min, n_max):
2 """
[…]
4 This function prints results directly based on the existence of a perfect square.
5 """
6 epsilon = 1e-9 # Tolerance for floating-point comparison
7 for n in range(n_min, n_max + 1):
8 L = a / n # Left-hand side of the equation
9 for x in range(n // a + 1, 10 * n): # Searching for x > n/a
10 t_min = max(1, (2 * n * x) // ((a * x - n) ** 2)) # Lower bound for t
11 for t in range(t_min, t_min + 100): # Iterate over possible t values in a reasonable range
12 q_square = t**2 * (a*x - n)**2 - 2 * n * t*x
13 if q_square >= 0:
14 q = int(q_square**0.5)
15 if q * q == q_square: # Ensure q is a perfect square
16 y = t * (a*x - n) - q
17 z = t * (a*x - n) + q
18 if y > 0 and z > 0:
19 R = 1/x + 1/y + 1/z
20 is_true = abs(L - R) < epsilon
21 print(f"For n = {n}, x = {x}, y = {y}, z = {z}, t = {t}: L = {L:.4f}, R = {R:.4f}, Conjecture holds: {is_true}")
22 break
23
24 # Example usage:
25 a = 4#(a=4 for the classic erdos_strauss conjecture,we can change de value of a>=4 depending on the conjecture we want to test)
26 n_min = 2
27 n_max = 150
28 #for a fixed value of a , we can vary the interval[n_min,n_max] to extend n.
29 print("Testing Second Formula:")
30 second_formula_erdos_straus(a, n_min, n_max)
Appendix : Python Code for Testing the classic Erdős-Straus Conjecture for the Values Not Proven by Mordell (Second Formula)
The following code tests the generalized Erdős-Straus conjecture using the second formula specifically for values of $n$ where Mordell’s method was inconclusive.
1 def second_formula_erdos_straus_mordell(a, n_min, n_max):
2 """
[…]
4 for values of n where Mordell’s method was inconclusive.
5 """
6 epsilon = 1e-9 # Tolerance for floating-point comparison
7 mordell_exceptions_mod_840 = {1, 121, 169, 289, 361, 529, 841, 961, 1009, 1129, 1369, 1681, 1849, 1969} # Corrected list of exceptional cases
8 captured_cases = 0 # Counter for validated cases
9 total_cases = 0 # Total exceptional cases in range
10
11 print("Mordell’s exceptional cases in the given range:")
12 for n in range(n_min, n_max + 1):
13 if n % 840 in mordell_exceptions_mod_840:
14 total_cases += 1
15 print(n, end=", ")
[…]
17
18 print("Testing Second Formula for Mordell’s Exceptional Cases:")
19 for n in range(n_min, n_max + 1):
20 if n % 840 not in mordell_exceptions_mod_840:
21 continue # Ignore values where Mordell already proved the conjecture
22
23 L = a / n # Left-hand side of the equation
24 solution_found = False
25
26 for x in range(n // a + 1, 10 * n): # Searching for x > n/a
27 t_min = max(1, (2 * n * x) // ((a * x - n) ** 2)) # Lower bound for t
28 for t in range(t_min, t_min + 100): # Iterate over possible t values in a reasonable range
29 q_square = t**2 * (a*x - n)**2 - 2 * n * t*x
30 if q_square >= 0:
31 q = int(q_square**0.5)
32 if q * q == q_square: # Ensure q is a perfect square
33 y = t * (a*x - n) - q
34 z = t * (a*x - n) + q
35 if y > 0 and z > 0:
36 R = 1/x + 1/y + 1/z
37 is_true = abs(L - R) < epsilon
38 captured_cases += 1
39 print(f"For n = {n} (Mordell case), x = {x}, y = {y}, z = {z}, t = {t}: "
40 f"L = {L:.4f}, R = {R:.4f}, Conjecture holds: {is_true}")
41 solution_found = True
42 break
43 if solution_found:
44 break
45
46 if total_cases > 0:
47 percentage_captured = (captured_cases / total_cases) * 100
[…]
49 else:
50 print("No exceptional cases found in the given range.")
51
52 # Example usage:
[…]
54 n_min = 2
55 n_max = 2000 # Increase to test larger cases
56
57 print("Testing Second Formula for Mordell’s Exceptional Cases:")
58 second_formula_erdos_straus_mordell(a, n_min, n_max)
Users can execute this script in Python to test numerical values and verify the conjecture for different cases.
Appendix: python code for percentages of n captures and recalcitrant cases
1
2
3 def second_formula_erdos_straus(a, n_min, n_max):
4 """
[…]
6 Stops searching for a given n as soon as one solution is found.
7 Computes the percentage of n captured and lists the uncaptured n values.
8 """
9 epsilon = 1e-9 # Tolerance for floating-point comparison
10 captured_n = 0 # Counter for the number of n with at least one solution
11 total_n = n_max - n_min + 1 # Total n values in the range
12 uncaptured_ns = [] # List to store n values that were not captured
13
14 for n in range(n_min, n_max + 1):
15 L = a / n # Left-hand side of the equation
16 found = False # Flag to check if a solution is found for this n
17
18 for x in range(n // a + 1, 100 * n): # Searching for x > n/a
19 t_min = max(1, (2 * n * x) // ((a * x - n) ** 2)) # Lower bound for t
20
21 for t in range(t_min, t_min + 100): # Iterate over possible t values
22 q_square = t**2 * (a * x - n)**2 - 2 * n * t * x
23 if q_square >= 0:
24 q = int(q_square**0.5)
25 if q * q == q_square: # Ensure q is a perfect square
26 y = t * (a * x - n) - q
27 z = t * (a * x - n) + q
28 if y > 0 and z > 0:
29 R = 1/x + 1/y + 1/z
30 is_true = abs(L - R) < epsilon
31 print(f"For n = {n}, x = {x}, y = {y}, z = {z}, t = {t}: L = {L:.4f}, R = {R:.4f}, Conjecture holds: {is_true}")
32
33 captured_n += 1 # Increment the counter since we found a solution
34 found = True
35 break # Stop searching for more solutions for this n
36
37 if found:
38 break # Stop searching for this n as we already found a solution
39
40 if not found:
41 uncaptured_ns.append(n) # Store n if no solution was found
42
43 # Compute and print the percentage of captured n
44 percentage_captured = (captured_n / total_n) * 100
[…]
46
47 # Print uncaptured n values
48 if uncaptured_ns:
[…]
50 print(uncaptured_ns)
51 else:
[…]
53
54 # Example usage:
[…]
56 n_min = 2
57 n_max = 100
58 print("Testing Second Formula:")
59 second_formula_erdos_straus(a, n_min, n_max)
Appendix: python code for finding perfect square manually (Case $n=577$)
1
2 import math
3
4 def find_perfect_squares(start=18592, end=20000):
5 results = []
6 for t in range(start, end + 1): # Iterate over t in a reasonable interval
7 delta = t * (9 * t - 167330) # Correction here with 9 * t
8 if delta > 0: # Ensure delta is positive
9 sqrt_delta = math.isqrt(delta) # Integer square root of delta
10 if sqrt_delta * sqrt_delta == delta: # Check if it’s a perfect square
11 results.append((t, delta, sqrt_delta))
12
13 return results
14
15 # Execute the code
16 start_value = 25000
17 end_value = 40000 # Define a reasonable upper bound here
18 perfect_squares = find_perfect_squares(start=start_value, end=end_value)
19
20 # Display results
21 for t, delta, square_root in perfect_squares:
22 print(f"t = {t}, delta = {delta}, racine_delta = {square_root}")
23
24 if not perfect_squares:
25 print(f"No values of t between {start_value} and {end_value} produce a perfect square.")